# Stochastic trade-offs and the emergence of diversification in E. coli evolution experiments

### v2

@@ Abstract
Laboratory experiments with bacterial colonies, under well-controlled conditions often lead to evolutionary diversification, where at least two ecotypes emerge from an initially monomorphic population. Empirical evidence suggests that such ”evolutionary branching” occurs stochastically, even under fixed and stable conditions. This stochastic nature is characterized by: (i) occurrence in a significant fraction, but not all, of experimental settings, (ii) emergence at widely varying times, and (iii) variable relative abundances of the resulting subpopulations across experiments. Theoretical approaches to understanding evolutionary branching under these conditions have been previously developed within the (deterministic) framework of ”adaptive dynamics.” Here, we advance the understanding of the stochastic nature of evolutionary outcomes by introducing the concept of ”stochastic trade-offs” as opposed to ”hard” ones. The key idea is that the stochasticity of mutations occurs in a high-dimensional trait space and this translates into variability that is constrained to a flexible tradeoff curve. By incorporating this additional source of stochasticity, we are able to account for the observed empirical variability and make predictions regarding the likelihood of evolutionary branching under different conditions. This approach effectively bridges the gap between theoretical predictions and experimental observations, providing insights into when and how evolutionary branching is more likely to occur in laboratory experiments.
@@ I Introduction
Understanding the origin and evolution of biological diversity is one of the central issues in evolutionary biology [1, 2, 3]. Historically, diversification was primarily understood in the context of allopatry, resulting from geographical isolation between subpopulations of a common ancestral species [4, 5, 6, 7, 8]. However, in recent years there has been a shift to the consideration of sympatric diversification —which occurs without spatial segregation— driven by mounting empirical evidence from both field and laboratory experiments [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] as well as by the consolidation of adaptive dynamics theory as a robust framework for addressing this type of eco-evolutionary problems [23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
Typically, this evolutionary process culminates in finding a (local) maximum of the fitness landscape, where the population stabilizes. However, sometimes it may result in the counterintuitive phenomenon of ”evolutionary branching.” This occurs when the endpoint of the gradient-ascending dynamics, which simultaneously modifies the fitness landscape itself, is a fitness minimum rather than a maximum. At this type of ”singular point,” the population can only increase its overall fitness by splitting into two sub-populations. Each sub-population can then separately ascend one side of the fitness landscape, resulting in diversification into two distinct lineages [23, 24, 27, 32, 33, 34].
@@
Prior theoretical analyses within the adaptive dynamics framework have demonstrated that simplified eco-evolutionary models can undergo evolutionary branching under the described conditions [31, 10]. However, these approaches are deterministic in nature, leading to the prediction of branching occurring in all or none of the realizations based on certain modeling assumptions or parameter values. Specifically, to replicate branching, previous studies relied on the arbitrary selection of a mathematical function to represent a ”trade-off” between the different aspects to be optimized, while alternative choices of the trde-off curve did not result in branching [31, 10].
@@
Trade-offs, often encapsulated by the idea that “you cannot be good at everything,” are fundamental in ecological and evolutionary modeling [39]. Typically, hard trade-offs, prevalent in modeling, reflect the notion that evolutionary pressures drive organisms toward a Pareto front [40]. Here, improvements in one task come with a cost to efficiency in others [41]. In our context, a specific trade-off curve has been commonly assumed between the abilities to metabolize glucose and acetate. However, selecting a precise trade-off curve is often the least justified aspect of mathematical models. While conceptually well-founded, experimental measurement of trade-offs presents significant challenges [42, 43]. Furthermore, as extensively debated in the literature, the choice of a specific trade-off curve is pivotal as it fundamentally dictates the occurrence of evolutionary branching [24, 44]. An elegant solution to circumvent this conceptual challenge within the adaptive dynamics framework was proposed by de Mazancourt and Dieckmann [44]. They devised a strategy to assess, for any given eco-evolutionary model, the robustness of the potential for evolutionary branching as the mathematical form of the trade-off function is altered (see below).
@@
Our main goal here is to provide a theoretical and computational framework to account for the observed variability in the outcomes of controlled evolutionary-branching experiments. We aim to achieve this by introducing the concept of “stochastic” trade-offs, contrasting with the conventional ”hard” trade-offs, within the adaptive dynamics framework. Here, stochastic trade-off implies that the population is not bound to evolve along a rigid, predefined curve (Pareto front) in phenotypic space, as is the case with hard trade-offs. Instead, it operates within a more flexible, extended region, allowing for greater exploration of phenotypic diversity (see below).
Let us remark that recent theoretical studies have reported variability in the emergence of diversification across replicates [45, 46]. However, these studies employ different ecological models, which deviate from the experimental setups, differ in the evolutionary algorithms, and focus on distinct aspects. In particular, in [45] the strict trade-off is replaced by an upper bound for the consumption rates of the different resources and the authors investigate the effect of clonal interference (i.e. strong-mutations [47]) and its role in fostering diversification. On the other hand, the authors of [46] analyze the interplay between selection and diversification by controlling the relative rates of purely fitness mutations and resource-strategy mutations, examining the various emerging regimes.
@@
In our study, we depart from previous research in several key aspects: (i) we utilize a biologically realistic ecological model fed with empirically fitted data to quantitatively describe the chemostat experiments [14, 15, 16] (although the concepts could also be extended to analyze serial-dilution experiments); (ii) we propose a straightforward extension of the adaptive dynamics framework by adding stochasticity into the trade-off that allows for a controlled inclusion of variability; (iii) we generalize our results for a large family of trade-off curves, and (iv) we formulate additional theoretical predictions about the emergence of diversification under varied experimental scenarios.
@@ II.1 Ecological dynamics
Observe first that all growth kinetics are described using Monod’s functions [51]. For example, $V_{g}\frac{G}{K_{g}+G}$ (resp. $V_{a}\frac{A}{K_{a}+A}$) represents the consumption of glucose (resp. acetate) at a maximum growth rate per capita $V_{g}$ (resp. $V_{a}$) and half-saturation constant $K_{g}$ (resp. $K_{a}$). The factor involving $C_{g}$ controls the repression of acetate consumption due to the presence of glucose, while the one with $C_{a}$ represents the inhibition of overall growth caused by acetate-induced acidification of the medium. $\gamma$ is the conversion constant for biomass production from resources, $\beta$ determines the fraction of consumed glucose excreted as acetate, $D$ is the chemostat’s dilution rate, and $G_{0}$ is the rate at which glucose is supplied to the chemostat. The corresponding numerical values of all these parameters are taken from experimental measurements [14, 15, 16] and are specified in Table 1.
Thus, in a nutshell, the population $N$ increases due to the consumption of glucose and/or acetate. However, consumption of acetate is inhibited due to glucose repression and overall growth is limited due to the acidification of the medium caused by acetate.
@@ II.2 Evolutionary dynamics
To complete the eco-evolutionary model, we must define the evolutionary part of the dynamics. This involves determining which ecological parameters (e.g., constants in Eqs. 1-3) are subject to change over time and which remain constant throughout the evolutionary process. Based on experimental findings [14, 15, 16], it is evident that the abilities to metabolize both glucose and acetate evolve over the course of the experiments. Therefore, parameters such as $V_{a,g}$, $K_{a,g}$, and $C_{g}$ are all potential candidates for evolution. For the sake of simplicity, we opt for $V_{g}$ and $V_{a}$—the maximum growth rates in glucose and acetate, respectively, represented collectively as $\textbf{V}=(V_{g},V_{a})$—to serve as the evolving parameters. These parameters are allowed to vary across generations, defining the metabolic strategy of each strain.
@@
At this point, a specific trade-off function, denoted as $V_{a}=h(V_{g})$, is typically introduced. This function reflects the biological fact that enhancing one metabolic strategy comes at the expense of efficiency in the other, so that both strategies cannot be altered independently [39]. This implies that the canonical equation (Eq.(5)) needs to be modified by an additional term to account for such a hard constraint as specified in Methods Sec.VI.1 Eq.(6). Analyzing the fixed points of this equation, along with the properties (first and second derivatives) of the modified fitness gradient at these points, one could determine whether the course of evolution (in the deterministic limit) leads to an evolutionarily stable population, or to the emergence of evolutionary branching.  [^1] Analytical solutions to this problem are often challenging, so one frequently resort to pairwise invasibility plots for insights (see [24, 63] for further details).
@@
As discussed in the introduction and explicitly demonstrated in the Results section, the previously-described adaptive-dynamics model coupled with a hard trade-off is deterministic. This means that it either exhibits branching for every model realization or it does not for any, depending on the specific shape of the considered trade-off curve. Consequently, it fails to elucidate the observed variability in the experiments such as the timing of diversification (branching) events and the proportion of the two emerging ecotypes over time. Let us remark that stochastic effects in the context of adaptive dynamics have already been considered in the literature. Wakano and Iwasa [53] (and more recently some of us [34]) have shown that finite-size effects in the population number may prevent the system from undergoing evolutionary branching, even though it is theoretically predicted to emerge at a deterministic level. However, these effects become negligible for the very large population sizes under consideration ($N\sim 10^{13}$).
@@
To address stochasticity, we adopt a different approach. As stated in the introduction we are interested in explaining properties of the evolutionary experiments such as the time at which a diversification (branching) event occurs and the fraction of the two emerging ecotypes at a certain time. Therefore, we need to build a framework capable of reproducing actual stochastic evolutionary trajectories that follow, on average, the behavior described by the canonical equation Eq. (6), but also allow for variability.
@@
Ideally, one would aim to construct a bottom-up approach by defining an individual-based model where each individual has its own associated phenotype, allowing for evolution through the combined processes of mutation and selection to occur [34, 26]. However, given the size of empirical populations, typically around $N\sim 10^{13}$, such a ”microscopic” description is currently computationally unfeasible. Therefore, as an alternative, we employ a framework adapted from [54] (see Methods Sec. VI.3), which represents an effective ”mesoscopic” description of the evolutionary trajectories. This approach does not directly deal with individual organisms but rather with ”groups” of them (subpopulations) that share the same phenotype, i.e. individuals are clustered in groups and one writes dynamical equations for the abundances of such groups. Essentially, this involves considering a collection of possible coexisting subpopulations, which necessitates (i) expanding the set of ecological equations to account for multiple subpopulations instead of just one, (ii) running the ecological dynamics for a sufficiently large time as to reach a steady state, (iii) introducing an additional equation for a new mutant subpopulation whose phenotype slightly differs from the subpopulation from which it derives, and (iv) iterating this eco-evolutionary process.
@@
Two remarks are in order before proceeding: First, let us recall that this framework, where the ecological dynamics reach stationarity before new mutations appear —as customarily assumed in adaptive dynamics [27]— corresponds to the strong-selection weak-mutation regime [47]. However, note that for very large bacterial populations such as those studied here, the number of mutations per unit time in the entire population might be high [47]. Thus, the key implicit assumption is that, even if generic mutations occur at a high rate, those directly affecting pathways involving the metabolism of acetate and glucose are much rarer, allowing us to assume a separation of timescales between the ecological and evolutionary dynamics.
@@
Second, to achieve the strict limit of perfect time-scale separation in computational analyses, one would need to run the ecological dynamics for an arbitrarily long time to ensure that a steady state has been reached. Since the simulation time is necessarily finite, what one actually analyzes is a limit of a small, but not infinitely small, mutation rate. This is why the algorithm accommodates a set of possible subpopulations, i.e., a distribution in phenotypic space, instead of a single population. In any case, one can perform analyses by progressively expanding the time of the ecological dynamics to verify that the resulting distributions become sharper as the time scale is enlarged.
@@ II.3 Stochastic trade-offs
Figure 1: Schematic representation of the concept of stochastic trade-off. a) Metabolic constraints of bacteria influence the outcome of evolutionary experiments. b) These constraints can be represented by a (topologically complex) manifold (gray shadowed surface), in the multidimensional space formed by all the relevant variables for evolution, which in our case are $V_{g}$, $V_{a}$ and the rest, $\vec{v}$. Starting from the same point, the different evolutionary trajectories corresponding to different realizations of the experiments can move freely in the high-dimensional constraint manifold (golden curves), possibly giving raise to diversification events (yellow star). Thus, the trajectories in the ($V_{g},V_{a}$) plane (yellow lines) are restricted to a certain zone (blue area), but can be different from each other. We model this constraint region, that is, the stochastic trade-off in the lower-dimensional manifold, by adding a temporarily-correlated noise on top of a certain fixed curve that stands as the average of possible trajectories (inset plot).
To account for the experimental variability, we propose introducing an additional source of stochasticity. Specifically, we suggest the concept of a stochastic trade-off, wherein the population is not confined to a fixed trade-off curve in phenotypic space but can instead wander within a broader region with fuzzy boundaries [55]. This concept can be rationalized as follows (see also Fig.1).
The underlying metabolic constraints of bacteria are indeed the same for every realization of the evolutionary experiments. However, such constraints involve complex relations between the many variables that are susceptible to evolution. In our case, we explicitly consider the evolution of $V_{g}$, $V_{a}$ but, in principle, there are many other variables, denoted generically as $\vec{v}$, that could be affected by overall (energetic/metabolic/…) constraints. These intricate relationships can be mathematically modeled as a multidimensional constraint manifold, akin to a high-dimensional “Pareto front” as depicted in Fig. 1 [40, 56]. In other words, the actual trade-off involves a high-dimensional manifold, rather than the two-dimensional one characterizing metabolic preferences.
@@
More specifically, we describe these constraints in the $(V_{g},V_{a})$-plane as a stochastic process (modelled as an Ornstein-Uhlenbeck process with amplitude $\sigma$ and temporal correlation $\tau$, as explicitly shown in Methods Sec.VI.2), added as an orthogonal perturbation to some central trade-off curve. This curve stands as the average of trajectories and can be chosen with large flexibility without affecting the conclusions (see Sec.VI.2). Depending on the amplitude and temporal correlations of the additional stochasticity, trajectories are constrained to wander at tunable distances of the central trade-off curve, allowing us to model the flexibility of the metabolic constraints in the $(V_{g},V_{a})$-plane and in turn the variability of the trajectories.
It is important to note that, although selection in the two-dimensional phenotypic plane tries to push trajectories away from the average curve in the direction of pure fitness growth, where both $V_{g}$ and $V_{a}$ increase, our model ”forces” the trajectories to stay close to the average curve. This serves as an effective way to model the constrains arising from the higher-dimensional manifold where mutations can have a positive, negative or neutral effect when projected onto the $(V_{g}-V_{a})$-plane.
In summary, the stochastic trade-off allows for the population to navigate within a flexible phenotypic region (in the $(V_{g},V_{a})$-plane) rather than being restricted to a fixed trade-off curve, effectively capturing the impact of metabolic constraints.

### v3

@@ Abstract
Laboratory experiments with bacterial colonies under well-controlled conditions often lead to evolutionary diversification, where at least two ecotypes emerge from an initially monomorphic population. Empirical evidence suggests that such an ”evolutionary branching” occurs stochastically, even under fixed and stable conditions. This stochasticity is characterized by: (i) occurrence of branching in a significant fraction, but not all, of experimental settings, (ii) emergence at widely varying times, and (iii) variable relative abundances of the resulting subpopulations across experiments. Theoretical approaches to understanding evolutionary branching under these conditions have been previously developed within the (deterministic) framework of ”adaptive dynamics.” Here, we advance the understanding of the stochastic nature of evolutionary outcomes by introducing the concept of ”stochastic trade-offs” as opposed to ”hard” ones. The key idea is that the stochasticity of mutations occurs in a high-dimensional trait space and this translates into variability that is constrained to a flexible tradeoff curve in a lower-dimensional space. By incorporating this additional source of stochasticity, we are able to account for the observed empirical variability and make predictions regarding the likelihood of evolutionary branching under different experimental conditions. This approach effectively bridges the gap between theoretical predictions and empirical observations, providing insights into when and how evolutionary branching is more likely to occur in laboratory experiments.
@@ I Introduction
Understanding the origin and evolution of biological diversity is one of the central issues in evolutionary biology [1, 2, 3]. Historically, diversification was primarily understood in the context of allopatry, resulting from geographical isolation between subpopulations of a common ancestral species [4, 5, 6, 7, 8]. However, in recent years there has been a shift to the consideration of sympatric diversification, which occurs without spatial segregation. This has been driven by mounting empirical evidence from both field and laboratory experiments [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] as well as by the consolidation of adaptive dynamics theory as a robust theoretical framework for addressing this type of eco-evolutionary problems [23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
Typically, this evolutionary process culminates in finding a (local) maximum of the fitness landscape, where the population stabilizes. However, sometimes it may result in the counterintuitive phenomenon of ”evolutionary branching.” This occurs when the endpoint of the gradient-ascending dynamics, which simultaneously modifies the fitness landscape itself, is a fitness minimum rather than a maximum [23, 24, 25, 26, 27]. At this type of “singular point”, the population can only increase its overall fitness by splitting into two sub-populations. Each sub-population can then separately ascend one side of the fitness landscape, resulting in diversification into two distinct lineages [23, 24, 27, 32, 33, 34].
@@
Prior theoretical analyses within the adaptive dynamics framework have demonstrated that simplified eco-evolutionary models can undergo evolutionary branching under the described conditions [31, 10]. However, these approaches are deterministic in nature, leading to the prediction of branching occurring in all or none of the realizations based on certain modeling assumptions or parameter values. Specifically, to replicate branching, previous studies relied on the arbitrary selection of a mathematical function to represent a ”trade-off” between the different metabolic choices to be optimized, while alternative choices of the trade-off curve did not result in branching [31, 10].
@@
Trade-offs, often encapsulated by the idea that “you cannot be good at everything,” are fundamental in ecological and evolutionary modeling [39]. Typically, hard trade-offs, prevalent in modeling, reflect the notion that evolutionary pressures drive organisms toward a Pareto front [40]. Here, improvements in one task come with a cost to efficiency in others [41]. In our context, a specific trade-off curve has been commonly assumed between the abilities to metabolize glucose and acetate. However, selecting a precise trade-off curve is often the least justified aspect of mathematical models. While conceptually well-founded, experimental measurement of trade-offs presents significant challenges [42, 43]. Furthermore, as extensively debated in the literature, the choice of a specific trade-off curve is pivotal as it fundamentally dictates the occurrence or absence of evolutionary branching [24, 44]. An elegant solution to circumvent this conceptual challenge within the adaptive dynamics framework was proposed by de Mazancourt and Dieckmann [44]. They devised a strategy to assess, for any given eco-evolutionary model, the robustness of the potential for evolutionary branching as the mathematical form of the trade-off function is altered (see below).
@@
Our main goal here is to provide a theoretical and computational framework to account for the observed variability in the outcomes of controlled evolutionary-branching experiments. We aim to achieve this by introducing the concept of “stochastic” trade-offs, contrasting with the conventional ”hard” trade-offs, within the adaptive dynamics framework. Here, stochastic trade-off implies that the population is not bound to evolve along a rigid, predefined curve (Pareto front) in phenotypic space, as is the case with hard trade-offs. Instead, it operates within a more flexible and extended region, allowing for greater exploration of phenotypic diversity. As we will show, by adjusting the width of this region, which reflects the flexibility of the underlying bacterial metabolic constraints, it is possible to effectively regulate the degree of variability.
Let us remark that very recent theoretical works have also reported variability in the emergence of diversification across replicates [45, 46]. In particular, in [45] the strict trade-off is replaced by an upper bound for the consumption rates of the different resources (i.e. the metabolic traits) and the authors investigate the effect of clonal interference (i.e. strong-mutations [47]) and its role in fostering diversification under varying mutation rates. Conversely, in [46], the authors examine the interplay between selection and diversification by controlling the relative rates of purely fitness mutations and resource-strategy mutations, exploring the resulting regimes. Thus, in both studies, variability in diversification is controlled by mutation rates. We also find that larger mutation rates (beyond the infinitesimally small mutation limit) enhance variability. However, in contrast to previous studies, we propose an alternative source of variability that is independent of changes in the mutation rate: a stochastic trade-off model, where variability arises from the presence of a high-dimensional trade-off manifold representing the underlying metabolic constraints between all evolution-susceptible variables. Specifically, in our approach the variability of metabolic traits in the high-dimensional manifold is captured by a stochastic evolution of phenotypic traits in a lower dimensional space (see below).
@@
Given the limited availability of experimental data, it is challenging to determine which mechanism accounts best for the observed variability, and it is also possible that multiple factors may contribute. However, we believe that the framework we introduce is especially well-suited for investigating this problem because: (i) we utilize a biologically realistic ecological model fed with empirically fitted data to quantitatively describe the chemostat experiments [14, 15, 16] (although the concepts could also be extended to analyze serial-dilution experiments); (ii) we propose a straightforward extension of the well-known framework of adaptive dynamics by adding stochasticity into the trade-off which allows for a controlled inclusion of variability; (iii) we generalize our results for a large family of trade-off curves, and (iv) we formulate additional theoretical predictions about the emergence of diversification under varied experimental scenarios.
@@ II.1 Ecological dynamics
Observe first that all growth kinetics are described using Monod’s functions [51]. For example, $V_{g}\frac{G}{K_{g}+G}$ (resp. $V_{a}\frac{A}{K_{a}+A}$) represents the consumption of glucose (resp. acetate) at a maximum growth rate per capita $V_{g}$ (resp. $V_{a}$) and half-saturation constant $K_{g}$ (resp. $K_{a}$). The factor involving $C_{g}$ controls the repression of acetate consumption due to the presence of glucose, while the one with $C_{a}$ represents the inhibition of overall growth caused by acetate-induced acidification of the medium. $\gamma$ is the conversion constant for biomass production from resources, $\beta$ determines the fraction of consumed glucose excreted as acetate, $D$ is the chemostat’s dilution rate, and $G_{0}$ is the rate at which glucose is supplied to the chemostat. The corresponding numerical values of all these parameters are taken from experimental measurements [14, 15, 16] and are specified in SM Table I.
Thus, in a nutshell, the population $N$ increases due to the consumption of glucose and/or acetate. However, consumption of acetate is inhibited due to catabolite repression and overall growth is limited due to the acidification of the medium caused by acetate.
@@ II.2 Evolutionary dynamics
To complete the eco-evolutionary model, we must define the evolutionary part of the dynamics. This involves determining which ecological parameters (e.g., constants in Eqs. 1-3) are subject to change over time and which remain constant throughout the evolutionary process. Based on experimental findings [14, 15, 16], it is evident that the abilities to metabolize both glucose and acetate evolve over the course of the experiments. Therefore, parameters such as $V_{a,g}$, $K_{a,g}$, and $C_{g}$ are all potential candidates for evolution. For the sake of simplicity, we opt for $V_{g}$ and $V_{a}$ —the maximum growth rates in glucose and acetate, respectively, represented collectively as $\textbf{V}=(V_{g},V_{a})$— to serve as evolving parameters. These parameters are allowed to vary across generations, defining the metabolic strategy of each strain.
@@
At this point, a specific trade-off function, denoted as $V_{a}=h(V_{g})$, is typically introduced. This function reflects the biological fact that enhancing one metabolic strategy comes at the expense of efficiency in the consumption of the other, so that both strategies cannot be altered independently [39]. This implies that the canonical equation (Eq.(5)) needs to be modified by an additional term to account for such a hard constraint as specified in Supplementary Material (SM) Eq.(1). Analyzing the fixed points of the resulting equation, along with the properties (first and second derivatives) of the modified fitness gradient at these points, one could determine whether the course of evolution (in the deterministic limit) leads to an evolutionarily stable population, or to the emergence of evolutionary branching  [^1] Analytical solutions to this problem are often challenging, so one frequently resort to pairwise invasibility plots for insights (see [24, 63] for further details)..
@@
As discussed in the introduction and explicitly demonstrated in the Results section, the previously-described adaptive-dynamics model coupled with a hard trade-off is deterministic. This means that it either exhibits branching for every model realization or it does not for any, depending on the specific shape of the considered trade-off curve. Consequently, it fails to elucidate the observed variability in the experiments such as the timing of diversification (branching) events and the proportion of the two emerging ecotypes over time. Let us remark that stochastic effects in the context of adaptive dynamics have already been considered in the literature. Wakano and Iwasa [53] (and more recently some of us [34]) have shown that finite-size effects in the population number may prevent the system from undergoing evolutionary branching, even though it is theoretically predicted to emerge at a deterministic level. However, these effects become negligible for the very large population sizes under consideration (here, $N\sim 10^{13}$).
@@
To address stochasticity, we adopt a different approach. As stated in the introduction we are interested in explaining properties of the evolutionary experiments such as the time at which a diversification (branching) event occurs and the fraction of the two emerging ecotypes at a certain time. Therefore, we need to build a framework capable of reproducing actual stochastic evolutionary trajectories that follow, on average, the behavior described by the canonical equation, but also allow for variability.
@@
Ideally, one would aim to construct a bottom-up approach by defining an individual-based model where each individual has its own associated phenotype, allowing for evolution through the combined processes of mutation and selection to occur [34, 26]. However, given the size of empirical populations, typically around $N\sim 10^{13}$, such a ”microscopic” description is currently computationally unfeasible. Therefore, as an alternative, we employ a framework adapted from [54] (see SM Sec. III), which represents an effective ”mesoscopic” description of the evolutionary trajectories. This approach does not directly deal with individual organisms but rather with ”groups” of them (subpopulations) that share the same phenotype, i.e. individuals are clustered in groups and one writes dynamical equations for the abundances of such groups. Essentially, this involves considering a collection of possible coexisting subpopulations, which necessitates (i) expanding the set of ecological equations to account for multiple subpopulations instead of just one, (ii) running the ecological dynamics for a sufficiently large time as to reach a steady state, (iii) introducing an additional equation for a new mutant subpopulation whose phenotype slightly differs from the subpopulation from which it derives, and (iv) iterating this eco-evolutionary process.
@@
Two remarks are in order before proceeding: First, let us recall that this framework, where the ecological dynamics reach stationarity before new mutations appear —as customarily assumed in adaptive dynamics [27]— corresponds to the strong-selection weak-mutation regime [47]. Note that for very large bacterial populations such as those studied here, the number of mutations per unit time in the entire population might be high [47]. Thus, the key implicit assumption is that, even if generic mutations occur at a high rate, those directly affecting pathways involving the metabolism of acetate and glucose are much rarer, allowing us to assume a large separation of timescales between the ecological and evolutionary dynamics. Experimental findings support this assumption, showing that all acetate-scavenging strains acquired their ability to metabolize acetate through mutations in a specific gene [16]. Therefore, although the overall mutation rate might be high, mutations impacting that particular gene are much less frequent.
@@
Second, to achieve the strict limit of perfect time-scale separation in computational analyses, one would need to run the ecological dynamics for an arbitrarily long time to ensure that a steady state has been reached. Since the simulation time is necessarily finite, what one actually analyzes is a limit of a small, but not infinitely small, mutation rate, ı.e., a weak (but not infinitely-weak) mutation limit. This is why the algorithm accommodates a set of possible subpopulations, i.e., a distribution in phenotypic space, instead of a single population. In any case, one can perform analyses by progressively expanding the time of the ecological dynamics to verify that the resulting distributions become sharper as the time scale is enlarged.
@@ II.3 Stochastic trade-offs
Figure 1: Schematic representation of the concept of stochastic trade-off. a) Metabolic constraints of bacteria influence the outcome of evolutionary experiments. b) These constraints can be represented by a (topologically complex) manifold (gray shadowed surface), in the multidimensional space formed by all the relevant variables for evolution, which in our case are $V_{g}$, $V_{a}$ and the rest, $v$. Starting from the same point, the different evolutionary trajectories corresponding to different realizations of the experiments can move freely in the high-dimensional constraint manifold (golden curves), possibly giving rise to diversification events if they reach a singular point (yellow star). Thus, the trajectories projected in the ($V_{g},V_{a}$) plane (yellow lines) are restricted to a certain zone (blue area), but can be different from each other. We model this constraint region, that is, the stochastic trade-off in the lower-dimensional manifold, by adding a temporarily-correlated noise on top of a certain fixed curve that stands as the average of possible trajectories (inset plot).
We here introduce an additional source of stochasticity to account for the experimental variability. Specifically, we suggest the concept of a stochastic trade-off, wherein the population is not confined to a fixed trade-off curve in phenotypic space but can instead wander within a broader region with fuzzy boundaries [55].
The concept of stochastic tradeoff can be rationalized as follows (see also Fig.1). The underlying metabolic constraints of bacteria are indeed the same for every realization of the evolutionary experiments. However, such constraints involve complex relations between the many variables that are susceptible to evolution. In our case, we explicitly consider the evolution of $V_{g}$, $V_{a}$ but, in principle, there are many other variables, denoted generically as $\vec{v}$, that could be affected by overall (energetic/metabolic/…) constraints. These intricate relationships can be mathematically modeled as a multidimensional constraint manifold, akin to a high-dimensional “Pareto front” as depicted in Fig. 1 [40, 56]. In other words, the actual trade-off involves a high-dimensional manifold, rather than the two-dimensional one characterizing metabolic preferences.
@@
More specifically, we describe these constraints in the $(V_{g},V_{a})$-plane as a stochastic process (modelled as an Ornstein-Uhlenbeck process with amplitude $\sigma$ and temporal correlation $\tau$, as explicitly shown in SM Sec. II), added as an orthogonal perturbation to some central trade-off curve. This curve stands as the average of trajectories and can be chosen with large flexibility without affecting the conclusions (see SM Sec. II). Depending on the amplitude and temporal correlations of the additional stochasticity, trajectories are constrained to wander at tunable distances of the central trade-off curve, allowing us to model the flexibility of the metabolic constraints in the $(V_{g},V_{a})$-plane and in turn the variability of the trajectories.
It is important to note that, although selection in the two-dimensional phenotypic plane tries to push trajectories away from the average curve in the direction of increasing fitness, where both $V_{g}$ and $V_{a}$ growth, our model ”forces” the trajectories to stay close to the average curve. This serves as an effective way to model the constrains arising from the higher-dimensional manifold where mutations can have a positive, negative or neutral effect when projected onto the $(V_{g}-V_{a})$-plane.
In summary, the stochastic trade-off allows for the population to navigate and randomly explore a broad phenotypic region (in the $(V_{g},V_{a})$-plane), effectively capturing the impact of metabolic constraints, rather than being restricted to a fixed trade-off curve.